Write-Right, a vertically integrated firm produces both paper and writing tablets. The demand for tables is given by
PT = 1.00 0.001Q
where Q is the quantity of tablets. The marginal cost of producing the paper necessary for tablet is
MCP = 0.20 + 0.001Q
It costs the firm $0.10 to make the paper into a writing tablet. If there is no external market for the paper, what transfer price should top management set for the paper?
1
Expert's answer
2020-11-25T12:21:01-0500

If MCP = 0.20 + 0.001Q and PT = 1.00 – 0.001Q, MCP = PT,

"0.20 + 0.001Q = 1.00 \u2013 0.001Q"

Q = 800,

So PT = 1.00 – 0.001Q = 0.2 is a transfer price.